OpenStudy (anonymous):

3. Solve: 0 = 6x2 − 10x − 4. (1 point) x = −1 over 3, x = −2 x = −1 over 3, x = 2 x = 1 over 3, x = −2 x = 1 over 3, x = 2

OpenStudy (anonymous):

6x^2 -10x -4 3x^2 -5x -2 factors of -6 that add up to 5 are -6 and 1 3x^2 -6x +x - 2 3x(x-2)+1(x-2) (3x+1) (x-2)=0 x=-1/3 or 2

OpenStudy (anonymous):

so no negatives?

OpenStudy (anonymous):

That's the hard part for me

OpenStudy (anonymous):

-1/3 and 2

OpenStudy (anonymous):

3x+1=0 3x= -1 x= -1/3 or x-2=0 x=0+2 x=2